The mathmatics department of a college has 14 male professors, 14 female proffesors, 7 male teaching assistants, and 14 female teaching assistants. If a person is selected at random from the group, find the probability that the selected person is a teaching assistant or a female.
Answer by nyc_function(2741)

You need the formula: P(A or B) = P(A) + P(B) - P(A and B),
where is P = probability.
Let A = selecting a teaching assistance
Let B = selecting a female
P(A or B) = P(A) + P(B) - P(A and B)
P(A or B) = (28/49) + (28/49) - (28/49*28/49)
P(A or B) = (4/7) + (4/7) - (4/7*4/7)
P(A or B) = (8/7) - (16/49)
P(A or B) = 40/49